OpenStudy (anonymous):

what is the range of f(x)=cos(x) y less than or equal to 0?

OpenStudy (rsadhvika):

range of cos(x) is [-1, 1]

OpenStudy (rsadhvika):

whats the exact question ?

OpenStudy (anonymous):

either true or false\[the range of f(x)=cos(x) is y \le0\] sorry I asked wrong

OpenStudy (rsadhvika):

false

OpenStudy (anonymous):

ok thank you can you help me with more please?

OpenStudy (rsadhvika):

im happy to help if i can, but it makes sense oly if u r also understanding them hmm soo,,,

OpenStudy (anonymous):

what is the y value of cos(x) when x=90?

OpenStudy (rsadhvika):

y = cos(x)

OpenStudy (rsadhvika):

put x = 90 y = cos(90)

OpenStudy (rsadhvika):

whats cos(90) ?

OpenStudy (anonymous):

its multiple choice a. -1 b.0 c.1/2 d.1

OpenStudy (rsadhvika):

to answer that, u need to know cos(90)

OpenStudy (rsadhvika):

use ur calculator, or google and see wat u get for cos(90)

OpenStudy (anonymous):

so its 0?

OpenStudy (rsadhvika):

cos(90) = 0 yes !

OpenStudy (anonymous):

what is the domain of f(x)=cos(x)? what is the y value of tan(x) when x=-180 the range of \[f(x)=\sin(x)is-1 \le y \le 1\] soo sorry about all the questions I really don't understand this section

OpenStudy (anonymous):

@rsadhvika please help